If a,b,c are selected from the set{0,1,2,3,4} then find the probability that the equation ax^2+bx+c=0 has infinite roots.

Grade:12

1 Answers

Umang goyal
6 years ago
Total cases = 2^n =2^5 =32If a equation has both roots as infinity it means a=b=C=0And now u can easily found probability..